Please help me solve this equation world problem
If a plane can travel 480 miles per hour with the wind and 440 miles per hour against the wind, find the speed of the plane in still air

Answer by checkley77(12844)

s+w=480
s-w=400 add,
--------------
2s=880
s=880/2
s=440 mph. speed of the plane in still air.
(480-440)=40 mph. speed of the wind.